OpenStudy (anonymous):

A van traveling down a slope with a uniform acceleration of 2.15 meters/second2 attains a speed of 20.00 meters/second after 7.00 seconds. What is the initial velocity of the van?

OpenStudy (amistre64):

if memory serves, the formula vf^2 = vi^2 + 2at

OpenStudy (souvik):

use it...\[v=u+a t\] v=final velocity u = initial velocity a=acceleration t=time

OpenStudy (amistre64):

lol, soo close

OpenStudy (amistre64):

i believe mine is for the change in distance, not the time

OpenStudy (souvik):

v^2=u^2+2ad

OpenStudy (amistre64):

yeah, that one :)

OpenStudy (anonymous):

thanks!